OpenStudy (anonymous):

\[\sin \Pi/6*\sin 2\Pi/6*\sin 4\Pi/6*\sin 5\Pi/6\] is ?

OpenStudy (experimentx):

180/6 = 30 ... I think they all have standard values at the multiples of 30??

OpenStudy (experimentx):

and finish your other problem first ...

OpenStudy (anonymous):

The answer is 3/16

OpenStudy (anonymous):

@experimentX

OpenStudy (experimentx):

Jeez ... sin(pi - x) = sin(x) use the property for 5x and 4x ... they all have standard values ... put those and get your answer

OpenStudy (anonymous):

@FoolForMath plzzz help

OpenStudy (experimentx):

what have you tried so far ...?

OpenStudy (anonymous):

yes not getting

OpenStudy (anonymous):

i did nt understand ur way

OpenStudy (anonymous):

@ParthKohli help

OpenStudy (experimentx):

sin(30)=?? sin(60)=?? sin(120) = ? sin(150) = ??

OpenStudy (experimentx):

if you have difficult evaluating 150 or 120 sin(150) = sin(180 - 150) = sin(30) similarly sin(120) = sin(60) you know these standard values ... if you don't know ... check google for trigonometric table ... and \( \pi \) = 180 degre

OpenStudy (anonymous):

ok @experimentX thanzzz

OpenStudy (anonymous):

thanzzzz

OpenStudy (experimentx):

yw